OpenStudy (anonymous):

How do I solve 2w-6(w+8)=-12

OpenStudy (anonymous):

I added 6 to both sides and did my multiplication and am stuck here, 2w^2+16w=-6

OpenStudy (anonymous):

multiply out first on the left

OpenStudy (anonymous):

\[2w-6(w+8)=-12\] means \[2\times w -6\times (w+8)=-12\]

OpenStudy (anonymous):

first step is not to add, it is to multiply on the left using the distributive property and get \[2w-6w-48=-12\]

OpenStudy (anonymous):

that is always the first step if you have parentheses, remove the parentheses using the distributive law second step is to combine like terms you get \[2w-6w-48=12\\ -4w-48=-12\]

OpenStudy (anonymous):

NOW it is time to add \(48\) to both sides \[-4w=36\] and finally divide both sides by \(-4\)and you are done

OpenStudy (the_fizicx99):

Wish I could explained, though I don't know how to do LaTex coding. Satellite's explanation is very detailed, so I know his would've been far better than mine. :)

OpenStudy (anonymous):

Wow, that was awesome. Thank you @satellite73, you're a gift sent from above! Haha :)

OpenStudy (anonymous):

yw, my pleasure

OpenStudy (anonymous):

Could you possibly help explain to me another question?

OpenStudy (anonymous):

I have been trying to solve -5-(11-2g)=-(g+8)+5

OpenStudy (anonymous):

\[-5-(11-2g)=-(g+8)+5 \] again distribute this time on both sides \[-5-11+2g=-g-8+5\] now combine like terms, this time that means the numbers \[-16+2g=-g-3\]

OpenStudy (anonymous):

add \(g\) to both sides to get them both on the left \[-16+3g=-3\]

OpenStudy (anonymous):

add \(16\) to both sides so all the numbers are on the right \[3g=13\] and finally divide by \(3\) to get \(g=\frac{13}{3}\)

OpenStudy (anonymous):

becoming more clear i hope